Discuss with Askiitians Tutors

CAN I GET THE FORMULA FOR THE MOMENT OF INERTIA OF A THICK WALLED SPHERICAL SHELL (HAVING AN INNER AND OUTER RADII) ABOUT ITS TANGENTIAL AXIS OF ROTATION

Sajeev Korappath
7 Years agoGrade

2 Answers

Arun
6 Years ago
For a solid sphere I = 2/5*M*a^2. If density = d (can’t see how to type rho!) M= 4/3*pi*a^3*d so I = 8/15*pi*a^5*d. If the inside hollow has radius r, it has I = 8/15*pi*r^5*d.
 
So I for the hollow sphere = 8/15*pi*(a^5-r^5)*d = 8/15*pi*(a-r)(a^4+a^3*r+a^2*r^2+a^r^3+r^4)*d
 
If a-r
 
Mass of shell = 4*pi*a^2*t*d so I = 2/3*M*a^2